Graph the Given Function and its inverse. f(x)

3. three-fourth x + 4= y
1
Expert's answer
2013-07-23T07:35:19-0400
3/4 x+4=y it's a straight line
To graph it we need two distinct points of it. We can pick arbitrary values of x, say x=0 and x=4:
y(0)=4
y(4)=3/4*4+4=7
So we have two points (0,4) and (4,7). Connecting them we get desired graph.
To find inverse we should solve given equation for x:
3/4x+4=y
3/4x=y-4
x=4/3(y-4)=4/3y-16/3
It's also a straight line. We can find its graph similarly.

To get more info about straight lines watch our video: